avi8tor:

suppose you flip a coin n times, and the probability of getting heads 15 times is .0148. what's n?

Hero:

@avi8tor are you familiar with the Binomial Probability Formula? The Binomial Probability Formula: \[P(x) = \large\dfrac{n!}{x!(n - x)!}^{p^xq^{(n - x)}}\] Where \(x\) represents the number of successes \(n\) represents the number of trials \(p\) represents the probability of success \(q\) represents the probability of failure \(\left(\text{In other words }, q = 1 - p\right)\) In this case \(x\) represents the number of times heads is flipped \((x = 15)\) \(n\) represents the total number of times the coin is flipped \(\left(\text{What we have to find}\right)\) \(p\) represents the probability of flipping heads in one event \( p = \left(\dfrac{1}{2}\right)\) \(q\) represents the probability of flipping tails in one event \(q = 1 - \left(\dfrac{1}{2}\right) = \dfrac{1}{2}\)

Hero:

You can find \(n\) by plugging all the values in to the formula and using trial and error for n values. We know this much: \(n \ge 15\)

Hero:

BTW, \(P(x)\) represents the probability of achieving \(x\) number of successes out of \(n\) events.

Ultrilliam:

@avi8tor Your question has been answered, log on to see it!